# Hybrid classical-quantum communication networks

### v1

@@ 3.6 Coming full circle: quantum technology for classical communications
Recurring themes throughout this review have emphasized the value of leveraging existing technology from classical lightwave communications for quantum communications as well—unsurprising given the maturity of optical networks and the importance of integrating quantum systems into the existing fiber-optic infrastructure for practical deployment. Nevertheless, continued advances in quantum information processing systems lead us to predict a future filled with technology transfer in the opposite direction as well, as tools motivated by and developed for quantum networking find new applications in classical contexts, auspiciously benefiting the lightwave communications field that originally supported it. In this picture of the future, acute needs of quantum networks provide the impetus for solutions which in turn address complementary needs in classical networking that have hitherto lacked sufficient economic drivers.
There are several areas where the development of new quantum networking technology could be of benefit to conventional networks. In particular, lower-loss and lower-noise optical components could enable reduced digital signal processing to correct for transmission impairments, leading to improved energy efficiency. Low-loss optical circuitry represents one such example immediately applicable to classical use. As highlighted in Sec. 3.2, conventional optical amplification is unavailable to quantum signals, as expressed concisely by the no-cloning theorem [254, 60]. Consequently, loss cannot be compensated via EDFAs in quantum communications, fundamentally limiting the distance over which quantum information can be transmitted. In this milieu, multi-dB insertion losses—e.g., from electro-optic modulators, switches, and pulse shapers—that are perhaps acceptable inconveniences in classical communications suddenly become critical roadblocks in the quantum regime. On the quantum computing front, companies like PsiQuantum [4] are already driving the development of low-loss photonic integrated circuits, and we predict quantum networking will increasingly motivate research in low-loss communications components, both integrated and discrete. Such technology would then readily translate to classical communication links, enabling energy and cost savings by removing amplification or lowering launch powers.
@@
Among the various low-loss optical components, one particularly promising technology for both classical and quantum communication is the Hollow Core Nested Antiresonant Nodeless Fiber (HC-NANF) [180]. HC-NANFs offer several significant advantages over conventional glass-core fibers, including reduced optical nonlinearity, ultra-low latency, ultra-wide bandwidths, and the potential of lower transmission loss. Notably, the first hollow core fiber (HCF) to achieve loss levels (0.174 dB/km) comparable to conventional telecom fibers in the C-band was reported in [111]. The reduced nonlinearity of HC-NANFs results in lower crosstalk between classical and quantum channels in quantum/classical coexistence networks. This was demonstrated in a recent experiment where DV-QKD was successfully conducted alongside 1.6 Tbps classical traffic over a HCF link [18]. Although HC-NANF was not specifically designed for quantum networking, its potential in future quantum Internet applications may justify the cost of replacing existing optical fibers.
@@
On the architecture side, quantum information’s incompatibility with traditional OE conversion could foreshadow a renaissance in all-optical signal processing and packet switching, a possibility speculated in Secs. 2.1.5 and 2.2.4. Efficient error correcting quantum memory along with better network synchronization might also enable advanced concepts of all-optical classical networking. Given the more stringent constraints on noise and power levels in quantum applications, we anticipate that quantum all-optical solutions should generally transfer to the classical domain where they could increase bandwidth, lower latency, and reduce energy use. Indeed, illustrations of such quantum-to-classical translation have already emerged; we highlight one example in the context of the QFP introduced briefly in Sec. 2.2.4. Based on electro-optic phase modulators and Fourier transform pulse shapers, the QFP enables the scalable synthesis of arbitrary quantum gates in frequency-bin encoding [146]—a particularly appropriate DoF for stable and parallel transmission in SMF [123, 146, 140, 134]. Although inspired by and focused on quantum information processing in the majority of its experimental demonstrations, the QFP at bottom is a mode transformer: it coherently mixes and interferes discrete frequency bins in a controllable fashion. This mode transformation is optically linear and entirely independent of the input’s statistical properties, making the QFP equally accepting of classical states as quantum ones. Recognizing the opportunities presented by the QFP’s flexibility, previous research has explored classical all-optical processing both theoretically [137] and experimentally [141] for the specific operations of frequency broadcasting (where data at one input frequency is copied to $N$ outputs) and cyclic hops (where data at multiple input frequencies are transferred to distinct outputs). Reconfigurable, background-free, and based on the theoretical guarantees of the QFP, this all-optical approach can in principle support arbitrary wavelength conversion operations in a single platform, highlighting by example the value of considering quantum and classical applications together as quantum network technology evolves; their intersection is ripe for new ideas and crosspollination that can advance both.
4 Outlooks
@@
As the backbone of the modern Internet, fiber-optic communication networks have profoundly shaped human society and civilization. To date, most optical communication networks operate in classical communication modes, where information carried by light is fundamentally distinguishable. The emerging field of quantum communication expands the realm of optical communication by leveraging the generation, transmission, and detection of quantum superposition states of photons. This development unlocks revolutionary applications such as distributed quantum computing and ultra-secure communications. In this context, the envisioned quantum Internet promises a broader scope than today’s classical Internet, supporting both quantum and classical communication modalities.
@@
The pursuit of a global quantum Internet will drive the development of cutting-edge quantum technologies, including entangled photon sources, quantum memories, and quantum repeaters. Simultaneously, advancements in existing optical technologies—such as low-noise optical receivers, ultra-low-loss optical storage, novel optical fibers, and related components—will push the boundaries of classical communication. These innovations may enable previously impractical modalities, such as all-optical networks, to become viable solutions. In the long term, the Internet of the future is likely to consist of heterogeneous communication networks designed to seamlessly support both classical and quantum communication, ensuring a robust, versatile, and transformative global communication infrastructure.

### v2

@@ 3.6 Coming full circle: quantum technology for classical communications
Recurring themes throughout this review have emphasized the value of leveraging existing technology from classical lightwave communications for quantum communications—unsurprising given the maturity of optical networks and the importance of integrating quantum systems into the existing fiber-optic infrastructure for practical deployment. Nevertheless, continued advances in quantum information processing systems lead us to predict a future filled with technology transfer in the opposite direction as well, as tools motivated by and developed for quantum networking find new applications in classical contexts, auspiciously benefiting the lightwave communications field that originally supported it. In this picture of the future, acute needs of quantum networks inspire solutions which in turn address complementary needs in classical networking that have hitherto lacked sufficient economic drivers.
There are several areas where the development of new quantum networking technology could be of benefit to conventional networks. In particular, lower-loss and lower-noise optical components could enable reduced DSP to correct for transmission impairments, leading to improved energy efficiency. Low-loss optical circuitry represents one such example immediately applicable to classical use. As highlighted in Sec. 3.2, conventional optical amplification is unavailable to quantum signals [76, 77]. Consequently, loss cannot be compensated via EDFAs in quantum communications, fundamentally limiting the distance over which quantum information can be transmitted. In this milieu, multi-dB insertion losses—e.g., from electro-optic modulators, switches, and pulse shapers—that are perhaps acceptable inconveniences in classical communications suddenly become critical roadblocks in the quantum regime. On the quantum computing front, companies like PsiQuantum [304] are already driving the development of low-loss photonic integrated circuits, and we predict quantum networking will increasingly motivate research in low-loss communications components, both integrated and discrete. Such technology would then readily translate to classical communication links, enabling energy and cost savings by removing amplification or lowering launch powers.
@@
Among the various low-loss optical components, one particularly promising technology for both classical and quantum communication is the hollow core nested antiresonant nodeless fiber (HC-NANF) [27]. HC-NANFs offer several significant advantages over conventional glass-core fibers, including reduced optical nonlinearity, ultralow latency, ultrawide bandwidths, and the potential of lower transmission loss. Notably, the first hollow core fiber (HCF) to achieve loss levels (0.174 dB/km) comparable to conventional telecom fibers in the C-band was reported in [305]. The reduced nonlinearity of HC-NANFs results in lower crosstalk between classical and quantum channels in quantum/classical coexistence networks. This was demonstrated in a recent experiment where DV-QKD was successfully conducted alongside 1.6 Tbps classical traffic over a HCF link [198]. Although HC-NANF was not specifically designed for quantum networking, its potential in future quantum Internet applications may justify the cost of replacing existing optical fibers.
Given the central role of entanglement in quantum networking, a particularly promising advancement in classical communication enabled by quantum technologies is entanglement assisted communication (EACOMM). By utilizing preshared entanglement, EACOMM can significantly enhance classical communication rates [306], as exemplified by superdense coding [307, 308, 309, 310]. However, such schemes still operate within the limits imposed by the Holevo–Schumacher–Westmoreland (HSW) bound  [311, 312, 313], which defines the ultimate capacity for classical information transmission over quantum channels. More recently, by harnessing phase correlations between entangled bosonic modes, EACOMM surpassing the HSW bound has been demonstrated experimentally [314], marking a potentially transformative shift in communication theory. Although the generation and distribution of entanglement between remote parties remain challenging and resource-intensive, one can envision future hybrid classical-quantum networks where entanglement is distributed during off-peak hours, allowing for boosted classical communication efficiency during high-traffic periods.
@@
On the architecture side, quantum information’s incompatibility with traditional OE conversion could foreshadow a renaissance in all-optical signal processing and packet switching, a possibility speculated in Secs. 2.1.5 and 2.2.4. Efficient error-correcting quantum memory along with better network synchronization might also enable advanced concepts of all-optical classical networking. Given the more stringent constraints on noise and power levels in quantum applications, we anticipate that quantum all-optical solutions should generally transfer to the classical domain where they could increase bandwidth, lower latency, and reduce energy use. Indeed, illustrations of such quantum-to-classical translation have already emerged; we highlight one example in the context of the QFP introduced briefly in Sec. 2.2.4. Based on electro-optic phase modulators and Fourier-transform pulse shapers, the QFP enables the scalable synthesis of arbitrary quantum gates in frequency-bin encoding [185]—a particularly appropriate DoF for stable and parallel transmission in SMF [184, 185, 187, 188].
Although inspired by and focused on quantum information processing in the majority of its experimental demonstrations, the QFP at bottom is a mode transformer: it coherently mixes and interferes discrete frequency bins in a controllable fashion. This mode transformation is optically linear and entirely independent of the input’s statistical properties, making the QFP equally accepting of classical states as quantum ones. Recognizing the opportunities presented by the QFP’s flexibility, previous research has explored classical all-optical processing both theoretically [315] and experimentally [61] for the specific operations of frequency broadcasting (where data at one input frequency is copied to $N$ outputs) and cyclic hops (where data at multiple input frequencies are transferred to distinct outputs). Reconfigurable, background-free, and based on the theoretical guarantees of the QFP, this all-optical approach can in principle support arbitrary wavelength conversion operations in a single platform, highlighting by example the value of considering quantum and classical applications together as quantum network technology evolves; their intersection is ripe for new ideas and crosspollination that can advance both.
4 Outlook
@@
As the backbone of the modern Internet, fiber-optic communication networks have profoundly shaped human society and civilization. To date, most optical communication networks operate in classical communication modes, where information carried by light is fundamentally distinguishable. The emerging field of quantum communication expands the realm of optical communication by leveraging the generation, transmission, and detection of quantum superposition states of photons. This development unlocks revolutionary applications such as distributed quantum computing and ultrasecure communications. In this context, the envisioned quantum Internet promises a broader scope than today’s classical Internet, supporting both quantum and classical communication modalities.
@@
The pursuit of a global quantum Internet will drive the development of cutting-edge quantum technologies, including entangled photon sources, quantum memories, and quantum repeaters. Simultaneously, advancements in existing optical technologies—such as low-noise optical receivers, ultralow-loss optical storage, novel optical fibers, and related components—will push the boundaries of classical communication. These innovations may enable previously impractical modalities, such as all-optical networks, to become viable solutions. In the long term, the Internet of the future is likely to consist of heterogeneous communication networks designed to seamlessly support both classical and quantum communication, ensuring a robust, versatile, and transformative global communication infrastructure.